# Maximum gap in complex Ginibre matrices

### v1

@@ 5. Proofs of Main Results
Using the definitions of $\xi^{z!}$ and $\mathbf{L}$ together with (40), we find for $\mathbf{L}$-almost all $z\in B$,
$c^{-1}n\rho_{n}(\sqrt{n}z)\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]-\mathbb{E}[\xi_{n}(\sqrt{n}B\setminus\sqrt{n}T_{n,z})]\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
$\quad=c^{-1}n\rho_{n}(\sqrt{n}z)\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}(\rho_{n,\sqrt{n}z}^{(2)}(w,v)-\rho_{n,\sqrt{n}z}(w)\rho_{n}(v))\,\mathrm{d}v\mathrm{d}w$
$\quad=c^{-1}n\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}\rho_{n}^{(3)}(\sqrt{n}z,w,v)-\rho_{n}^{(2)}(\sqrt{n}z,w)\rho_{n}(v)\,\mathrm{d}v\mathrm{d}w.$
@@
$c^{-1}n\left|\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}\rho_{n}^{(3)}(\sqrt{n}z,w,v)-\rho_{n}^{(2)}(\sqrt{n}z,w)\rho_{n}(v)\,\mathrm{d}v\mathrm{d}w\right|$
$\quad\leq c^{-1}3^{5/2}\pi|B|n^{2}\sup_{z\in\sqrt{n}B}r_{n}(z)^{2}\sup_{\begin{subarray}{c}v,w\in\sqrt{n}B\\ |v-w|\geq\log n-r_{n}(\sqrt{n}z)\end{subarray}}|K_{n}(w,v)|$
(41) $\quad\leq c^{-1}3^{5/2}\pi|B|n^{2}2\sqrt{\log n}\sup_{\begin{subarray}{c}v,w\in\sqrt{n}B\\ |v-w|\geq\log n-2\sqrt{\log n}\end{subarray}}e^{-|w-v|^{2}/2}+O(e^{-cn})\leq\frac{2\cdot 3^{5/2}\pi|B|}{cn}+O(e^{-cn}).$
Hence, we obtain by adding the contributions from (35), (39) and (41),
(42) $\mathbb{E}[(\xi_{n}\Delta\xi_{n}^{z,\varnothing})(\sqrt{n}S_{n,z})]\leq 2|B|\Big{(}\frac{\pi}{n^{2}}+\frac{2\sqrt{\log n}+3^{5/2}\pi}{cn}\Big{)}+O(e^{-cn}),\quad z\in B.$
@@
By Lemma 8 there exists $\delta>0$ such that $B_{r_{n}}(w)\subset B_{(1-\delta)\sqrt{n}}(0)$ for all $w\in B$. Hence, the previous sum is bounded by
@@
from Lemma 8, we find by a straightforward volume argument (considering how many balls with radius $r_{n}(w)/2$ fit into a ball with radius $2r_{n}(v)$) that $\sum_{w\in\omega\cap B_{r_{n}}(v)}\mathds{1}\{\omega(B_{r_{n}}(w))=0\}\leq 16$ for all locally finite $\omega\in\mathbf{N}$ and $n$ large enough. Inserting this estimate into (LABEL:e:bound29) gives that (32) is almost surely bounded by
@@
Hence, by (42) (with $B$ replaced by $B_{(1-\delta)\sqrt{n}}(0)$) gives the bound
$32|B_{1-\delta}(0)|\Big{(}\frac{\pi}{n^{2}}+\frac{2\sqrt{\log n}+3^{5/2}\pi}{cn}\Big{)}+O(e^{-cn}),\quad z\in B.$
This shows the assertion. ∎
@@
as $n\to\infty$. Using the definition of $\Xi_{c,n}$, that $\mathbf{L}$ is the intensity measure for $\Xi_{c,n}$, and $\rho_{n}(z)\leq\pi^{-1}$ for all $z\in B$, we have
@@
for some $\delta\in(0,1)$, where in the last inequality we used that by [40, Theorem 1], there exists a coupling $(\xi^{\prime},\xi^{\prime\prime})$ of $\xi_{n}^{z!}$ and $\xi_{n}$ such that $\xi^{\prime}\subset\xi^{\prime\prime}$ and $|\xi^{\prime\prime}\setminus\xi^{\prime}|\leq 1$. By Lemma 7 with $s_{n}=(\sqrt{2}+\epsilon)\sqrt[4]{\log n}$ we have
@@
Hence, (44) tends to 00 as $n\to\infty$.
@@
Fix $c>0$. By (5), we have $(\sqrt{2}-\epsilon)\sqrt[4]{\log n}\leq r_{c,n}(z)$ for $n$ large enough, uniformly in $z\in\sqrt{n}B$. Therefore,
$\mathbb{P}\Big{(}\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{z\}}|z-w|\leq(\sqrt{2}-\varepsilon)\sqrt[4]{\log n}\Big{)}$ $\leq\mathbb{P}\Big{(}\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{z\}}|z-w|\leq r_{c,n}(z)\Big{)}$
$=\mathbb{P}(\Xi_{c,n}\cap B=\varnothing),$
which tends to $\mathbb{P}(\zeta_{c}\cap B=\varnothing)=\exp(-c|B|)$ as $n\to\infty$ by Theorem 1, since convergence in KR distance implies convergence in distribution for locally finite point processes (see [17, Proposition 2.1]). Since $c$ was arbitrary, this upper bound can be made arbitrarily small by taking $c$ sufficiently large. ∎
Acknowledgments. MO was supported by the NWO Gravitation project NETWORKS under grant agreement no. 024.002.003.

### v2

@@ 5. Proofs of Main Results
Using the definitions of $\xi^{z!}$ and $L_{n}$ together with (48), we find for $L_{n}$-almost all $z\in B$,
$\kappa^{-1}n\rho_{n}(\sqrt{n}z)\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]-\mathbb{E}[\xi_{n}(\sqrt{n}B\setminus\sqrt{n}T_{n,z})]\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
$\quad=\kappa^{-1}n\rho_{n}(\sqrt{n}z)\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}(\rho_{n,\sqrt{n}z}^{(2)}(w,v)-\rho_{n,\sqrt{n}z}(w)\rho_{n}(v))\,\mathrm{d}v\,\mathrm{d}w$
$\quad=\kappa^{-1}n\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}\rho_{n}^{(3)}(\sqrt{n}z,w,v)-\rho_{n}^{(2)}(\sqrt{n}z,w)\rho_{n}(v)\,\mathrm{d}v\,\mathrm{d}w.$
@@
$\kappa^{-1}n\left|\int_{B_{r_{n}}(\sqrt{n}z)}\int_{\sqrt{n}S_{n,z}}\rho_{n}^{(3)}(\sqrt{n}z,w,v)-\rho_{n}^{(2)}(\sqrt{n}z,w)\rho_{n}(v)\,\mathrm{d}v\,\mathrm{d}w\right|$
$\quad\leq\kappa^{-1}3^{5/2}\pi|B|n^{2}\sup_{z\in\sqrt{n}B}r_{n}(z)^{2}\sup_{\begin{subarray}{c}v,w\in\sqrt{n}B\\ |v-w|\geq\log n-r_{n}(\sqrt{n}z)\end{subarray}}|K_{n}(w,v)|$
(49) $\quad\leq\kappa^{-1}3^{5/2}\pi|B|n^{2}2\sqrt{\log n}\sup_{\begin{subarray}{c}v,w\in\sqrt{n}B\\ |v-w|\geq\log n-2\sqrt{\log n}\end{subarray}}e^{-|w-v|^{2}/2}+O(e^{-cn})\leq\frac{2\cdot 3^{5/2}\pi|B|}{cn}+O(e^{-cn}).$
Hence, we obtain by adding the contributions from (43), (47) and (49) that for all $z\in B$,
(50) $\mathbb{E}[(\xi_{n}\Delta\xi_{n}^{z,\varnothing})(\sqrt{n}S_{n,z})]\leq 2|B|\Big(\frac{\pi}{n^{2}}+\frac{2\sqrt{\log n}+3^{5/2}\pi}{cn}\Big)+O(e^{-cn}).$
This completes the bound of the expectation of the right-hand side of (39).
@@
By Lemma 13 there exists $\delta>0$ such that $B_{r_{n}}(w)\subset B_{(1-\delta)\sqrt{n}}(0)$ for all $w\in B$ and $n\geq C(s,\kappa)$. Hence, the previous sum is bounded by
@@
from Lemma 13, we find by a straightforward volume argument (considering how many balls with radius $r_{n}(w)/2$ fit into a ball with radius $2r_{n}(v)$) that
$\sum_{w\in\omega\cap B_{r_{n}}(v)}\mathds{1}\{\omega(B_{r_{n}}(w))=0\}\leq 16$
for all locally finite $\omega\in\mathbf{N}$ and $n$ large enough. Inserting this estimate into (LABEL:e:bound29) gives that (40) is almost surely bounded by
@@
Hence, by (50) (with $B$ replaced by $B_{(1-\delta)\sqrt{n}}(0)$), we have for all $z\in B$ that (40) is bounded by
$32|B_{1-\delta}(0)|\Big(\frac{\pi}{n^{2}}+\frac{2\sqrt{\log n}+3^{5/2}\pi}{cn}\Big)+O(e^{-cn}).$
This completes the bound for (40), and hence completes the proof. ∎
@@
as $n\to\infty$. Using the definition of $\Xi_{n}$, that $L_{n}$ is the intensity measure for $\Xi_{n}$, and $\rho_{n}(z)\leq\pi^{-1}$ for all $z\in B$, we have
@@
where in the last inequality we used the coupling of of $\xi_{n}^{z!}$ and $\xi_{n}$ indicated in Section 3.3. By Lemma 12 with $s_{n}=(\sqrt{2}+\epsilon)\sqrt[4]{\log n}$ we have for $n\geq C(\epsilon,B)$ that
@@
Hence, (52) tends to 0 as $n\to\infty$.
@@
Fix $\kappa>0$. By (6), we have $(\sqrt{2}-\epsilon)\sqrt[4]{\log n}\leq r_{n}(z)$ for $n\geq C(\epsilon,\kappa,B)$. Therefore,
$\mathbb{P}\Big(\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{z\}}|z-w|\leq(\sqrt{2}-\varepsilon)\sqrt[4]{\log n}\Big)$ $\leq\mathbb{P}\Big(\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{z\}}|z-w|\leq r_{n}(z)\Big)$
$=\mathbb{P}(\Xi_{n}\cap B=\varnothing),$
which tends to $\mathbb{P}(\zeta_{\kappa}\cap B=\varnothing)=\exp(-\kappa|B|)$ as $n\to\infty$ by Theorem 1, since convergence in KR distance implies convergence in distribution for locally finite point processes (see [18, Proposition 2.1]). Since $\kappa$ was arbitrary, this upper bound can be made arbitrarily small by taking $\kappa$ sufficiently large. ∎
Acknowledgments. PL was partially supported by NSF grant DMS-2450004. MO was supported by the NWO Gravitation project NETWORKS under grant agreement no. 024.002.003. The authors thank the anonymous referees for their many helpful comments.